OpenStudy (anonymous):

At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path.The bicyclist heading north is riding 5 km/hour faster than the bicyclist heading south. At 10:45, they are 47.25 km apart. Find the two bicyclists’ rates.

OpenStudy (anonymous):

(Scratching head)

OpenStudy (shinalcantara):

let x be the speed of bicyclist A (heading north) y be the speed of bicyclist B (heading south) since A is 5km/hr faster than B x = y + 5 -----equation 1 time = 9 to 10:45 = 1.75hours since the distance is 47.25km then, x(1.75) + y(1.75) = 47.25 ---equation 2 substitute equation 1 to 2 (y+5)(1.75) + 1.75y = 47.25 1.75y +8.75 + 1.75y = 47.25 3.5y = 47.25 - 8.75 3.5y = 38.5 y = 11km/hr substitute to equation 1 x = y + 5 x = 11 + 5 x = 16km/hr